Charles has twice as much money as Dave. Ethan has 3 times as much money as
Charles. If Ethan has $36 more than Charles, what was the total amount of
money the three boys had?

Answer by Edwin McCravy(20064)

Instead of doing yours for you, I'll do one exactly like yours with
different numbers.

Charles has three times as much money as Dave. Ethan has 4 times as much
money as Charles. If Ethan has $45 more than Charles, what was the total
amount of money the three boys had?


Charles has 3 times as much money as Dave.
C = 3D

Ethan has 4 times as much money as Charles.
E = 4C

If Ethan has $45 more than Charles,
E = C + 45

Since 4C and C + 45 both equal to E, they are equal to each other.

So we set them equal:

4C = C + 45
3C = 45
 C = 15 

Substitute 15 for C in

 C = 3D
15 = 3D
 5 = D

Substitute 15 for C in

E = 4C

E = 4(15)

E = 60

So Charles had $15, Dave had $5 and Ethan had $60.

what was the total amount of money the three boys had?

Together they had $15 + $5 + $60 = $80

Now do yours.

Edwin